ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

10. Partial Differential Equations

1991 MS Classification: 35 Palagachev, Dian K., University of Chemical Technology and Metallurgy, Sofia, Bulgaria Oblique derivative problem for uniformly elliptic operators with VMO coefficients and applications Strong solvability in Sobolev spaces W^{2,p}(Ω) will be discussed for the regular oblique derivative problem for second order uniformly elliptic operators with V MO-principal coefficients. The results will be applied to the degenerate (tangential) oblique derivative problem in the case of neutral vector field prescribing the boundary operator. The results are obtained in collaboration with Antonino Maugeri.